# Disk patterns, quasi-duality and the uniform bounded diameter conjecture

### v1

@@ 1. Introduction
Let $\widetilde{G}\subseteq\operatorname{PSL}_{2}(\mathbb{C})$ be a geometrically finite Kleinian group with connected limit set, and $M\coloneqq\widetilde{G}\backslash(\mathbb{H}^{3}\cup\Omega(\widetilde{G}))$ be the corresponding Kleinian 3-manifold, where $\Omega(\widetilde{G})$ is the domain of discontinuity of $\widetilde{G}$. The quasiconformal deformation space of $\widetilde{G}$ can be naturally identified with the Teichmüller space $\operatorname{Teich}(\partial M)$. The covering of $\operatorname{Int}(M)$ corresponding to $\partial M$ is a (potentially disconnected) quasifuchsian manifold whose conformal boundary is the union of $X\in\operatorname{Teich}(\partial M)$ and its skinning surface $\sigma_{M}(X)\in\operatorname{Teich}(\overline{\partial M})$. This defines a map between Teichmüller spaces
@@
Suppose that $M$ is acylindrical, or equivalently, the limit set $\Lambda(\widetilde{G})$ is homeomorphic to a round Schottky set, i.e., the complement of infinitely many disjoint round open disks in $\widehat{\mathbb{C}}$ (see Figure 1(c) and 3(b)). Thurston’s Bounded Image Theorem (see §2.5 and [Thu86]), which is a crucial step in Thurston’s hyperbolization theorem, states that the image of $\sigma_{M}$ has compact closure. Thus,
@@
Let $G$ be a discrete group generated by reflection along circles in $\hat{\mathbb{C}}$ with connected limit set, and let $\operatorname{QC}(G)$ be the quasiconformal deformation space of $G$. Let $\widetilde{G}\lhd G$ be the index 2 subgroup consisting of orientation preserving elements. Then $\widetilde{G}$ is a geometrically finite Kleinian group. Let $M\coloneqq\widetilde{G}\backslash(\mathbb{H}^{3}\cup\Omega(\widetilde{G}))$ be the corresponding Kleinian orbifold. Then
@@
  • the reflection group $G$ can contain parabolic elements, so $\widetilde{G}\backslash\mathbb{H}^{3}$ can contain cusps, thus the injectivity radius can be 00;
@@
  • for every $N\geq 3$, one can construct a sequence of acylindrical reflection groups with totally geodesic convex core boundary and topological complexity $N$, the depth of whose collar about the convex core boundary goes to 00;
@@ 1.1. Disk pattern and reflection group
Let $\mathcal{G}$ be a connected simple plane graph with vertex and edge sets $\mathcal{V}$ and $\mathcal{E}$. Let $\omega:\mathcal{E}\longrightarrow\{\frac{\pi}{n}\colon n\in\mathbb{N}_{\geq 2}\}\cup\{0\}$ be some weight function on its edge set. We call such an edge-weighted graph $(\mathcal{G},\omega)$ a Coxeter graph. A disk pattern with combinatorics $(\mathcal{G},\omega)$ is essentially a collection of disks $\mathcal{P}=\{D_{v},v\in\mathcal{V}\}$ whose intersection pattern and angles are described by the graph $\mathcal{G}$ and the weight function $\omega$ respectively (see Definitions 2.1 and 2.2 for the subtleties involving a parabolic face).
@@
Hence in our setting, the skinning map can be explicitly defined as follows. Given a disk pattern $\mathcal{P}$ associated to the reflection group, the input of the skinning map is represented by a collection of polygons $\{\Pi_{F,\mathcal{P}}\}_{F}$. For each hyperbolic face $F$, the corresponding component of the skinning image is revealed by removing the disks for vertices in the complement of $F$, and represented by another polygon $\Pi_{F,\mathcal{P}}^{-}$. See Figure 1.1 for an example illustrating the correspondence between reflection groups and circle patterns, as well as this explicit presentation of the skinning map.
(a) The Coxeter graph, where $n$ on an edge means weight $\pi/n$. Note that all but one hyperbolic face is triangular.
(b) The disk pattern. Removing the light grey disk reveals the image of the skinning map, which is the hyperbolic polygon marked in red.
(c) The limit set of the corresponding acylindrical reflection group. It is homeomorphic to a Sierpinski carpet.
Figure 1.1. An example of the correspondence between reflection groups and disk patterns.
@@ 1.2. Discrete extremal lengths / widths
We remark that the uniform lower bound in Equation (1.1) is the key in our argument to obtain a uniform upper bound on the skinning diameter. It is a discrete analogue of the reciprocal condition in [Raj17, NR22] (see §1.4.3 for more discussions).
@@
We remark that as one varies reflection groups with the same topological complexity, the hyperbolic polygons in the skinning image can become degenerate. Our Theorem 1.2, in particular Equation (1.3), states that different hyperbolic polygons in the skinning image must degenerate in the same way.
@@ 1.4.2. Circle packings and discrete extremal lengths
Circle packings (or, more generally, disk patterns), their deformation spaces, and rigidity problems have been studied in [RS87, Sch91, He99, RHD07, HL13, HL17]. More generally, these problems have been studied for circle packings on complex projective surfaces [Thu22, KMT03, KMT06, Lam21, BW23]. Circle packings and disk patterns have also long been employed to study geometric structures and their deformation spaces [Bro85, Bro86, FS97]. Recently, [LZ23] has explored the connections between the skinning map, renormalization, and circle packings (see also [LZ24]). Our main result suggests that it may be possible to extend the uniform contraction property of the renormalization operator, as discussed in [LZ23], to cases with non-fixed combinatorics.
@@
Quasi-duality has been applied to prove uniformization theorems of metric surfaces [Raj17, RR19, RRR21, EBPC22, Iko22, MW24]. Notably, in [Raj17], it is shown that a metric space homeomorphic to $\mathbb{R}^{2}$ is quasiconformally homeomorphic to an open domain in $\mathbb{C}$ if and only if it is reciprocal, which in particular requires uniform quasi-duality for all quadrilaterals in the space. This result is generalized in [NR22].
@@ Acknowledgement
The authors thank Y. Minksy for asking the question and suggesting the connection between uniform diameter bound and circle packings.
@@ 2.1. Realizable Coxeter graphs
Let $\mathcal{G}$ be a connected simple plane graph, and let $\mathcal{V}$ and $\mathcal{E}$ be the set of vertices and edges of $\mathcal{G}$ respectively. Let $\omega:\mathcal{E}\longrightarrow[0,\pi/2]$ be a weight function on the set of edges. We call such an edge-weighted graph a Coxeter graph.
@@
It is easy to see that if $F$ is parabolic, then it is either a triangle or a quadrilateral. It is also easy to see that if $F$ is hyperbolic with $n$ sides, then the weights on the edges add up to $<(n-2)\pi$.
For a more uniform presentation, we always assume that in the Coxeter graph $(\mathcal{G},\omega)$, there does not exist a pair of adjacent triangular parabolic faces sharing an edge with weight 00. If there does exist such a pair, it is easy to see that the four edges other than the one shared by the two triangles all have weight $\pi/2$. We can remove the common edge, and combine the two parabolic faces into a single quadrilateral parabolic face. This in fact does not affect the combinatorics of the disk patterns encoded by the graph, see Definitions 2.1 and 2.2, as well as the discussion between them.
@@
¯$\overline{\mathcal{E}}=\mathcal{E}\cup\bigcup_{F\in\mathcal{F}_{pq}}\{e^{F}_{13},e^{F}_{24}\}.$
The weight function satisfies $\overline{\omega}=\omega$ on $\mathcal{E}$, and $\overline{\omega}\equiv 0$ on $\¯\overline{\mathcal{E}}\backslash\mathcal{E}$.
@@
A disk pattern with Coxeter graph $(\mathcal{G},\omega)$ is a collection of closed round disks $\mathcal{P}\coloneqq\{D_{v},v\in\mathcal{V}\}$ so that
@@
(Triangulation of a general graph) Suppose now $(\mathcal{G},\omega)$ has at least one hyperbolic face. For each hyperbolic face $F$, add a vertex $v_{F}$ in its interior and connect it to all vertices on $\partial F$. We also assign weight 00 to all the new edges. Note that in this new graph, each hyperbolic face with $n$-sides is divided into $n$ hyperbolic triangles. Now for each new hyperbolic triangle, add an additional vertex and connect it to the three vertices of the triangle. We assign weight $\pi/2$ to these new edges. It is not hard to see that the new Coxeter graph has at least $7$ vertices, and has no hyperbolic face. Conditions (A) and (B) remain true. We can thus apply Step 2 and Step 3 to conclude that the new graph is realizable. Removing the additional disks, we conclude that the original graph is realizable.
@@ 2.2. Reflection groups associated to a Coxeter graph
Note that the infinite ends of $\mathcal{H}_{P}$ extend to the sphere at infinity $\hat{\mathbb{C}}$ as interstices. In particular, we conclude that $\mathcal{H}_{P}$ is nonempty and in fact has nonempty interior. By Poincaré’s polyhedron theorem, $G$ is a discrete subgroup of $\operatorname{Aut}^{\pm}(\hat{\mathbb{C}})$, and $\mathcal{H}_{\mathcal{P}}$ is a fundamental domain for the action of $G$ on $\mathbb{H}^{3}$. The group $G$ contains an index-2 subgroup of orientation-preserving elements, which we denote by $\widetilde{G}$. It follows from the construction of a fundamental domain above that $\widetilde{G}$ is geometrically finite, i.e. the action of $\widetilde{G}$ on $\mathbb{H}^{3}$ has a finite-sided fundamental polyhedron.
Let $\Lambda$ and $\Omega$ be the limit set and domain of discontinuity of $\widetilde{G}$ respectively. Let $\Pi_{F}$ be an interstice and suppose $G_{F}$ is the subgroup of $G$ generated by vertices $v\in\partial F$. Define $\Omega_{F}=\bigcup_{g\in G_{F}}g\cdot\overline{\Pi_{F}}$.
@@
For Part 1, it suffices to show that for any hyperbolic face $F$, the subgroup $G_{F}$ is nonelementary. Indeed, since $\mathcal{H}_{\mathcal{P}}$ gives a fundamental domain, $G_{F}\cdot\overline{\Pi_{F}}$ tiles $\Omega_{F}$. If $G_{F}$ is elementary, then ^$\hat{\mathbb{C}}-\Omega_{F}$ consists of at most $2$ points. On the other hand, if we set $\widetilde{G}_{F}$ to be the subgroup of orientation-preserving elements in $G_{F}$, then $\widetilde{G}_{F}\backslash\Omega_{F}$ is hyperbolic, so $\Omega_{F}$ admits a hyperbolic metric as well. This is a contradiction.
@@ 2.3. Moduli spaces and reflection groups
A quasiconformal deformation of $G$ is a discrete and faithful representation $\xi:G\longrightarrow\operatorname{Aut}^{\pm}(\hat{\mathbb{C}})$ that preserves parabolics, induced by a quasiconformal map $f:\hat{\mathbb{C}}\longrightarrow\hat{\mathbb{C}}$ (i.e. $\xi(g)=f\circ g\circ f^{-1}$ for any $g\in G$).
@@
where $\xi\sim\xi^{\prime}$ if they are conjugates of each other by a Möbius transformation.
@@
Consider the Kleinian 3-orbifold $M\coloneqq\widetilde{G}\backslash(\mathbb{H}^{3}\cup\Omega)$. Its boundary $\partial M=\bigcup_{F}X_{F}$ has a connected component for each hyperbolic face $F$ of $\mathcal{G}$. Conformally, $X_{F}\cong\widetilde{G}_{F}\backslash\Omega_{F}$. In fact, $X_{F}$ can be constructed as the double of $\Pi_{F}$, with punctures or cone points of order $n$ if the corresponding edges have weight 00 or $\pi/n$. The surface $X_{F}$ has an anti-conformal involution $r=r_{F}$ given by exchanging the two copies of $\Pi_{F}$.
@@ 2.4. Limit sets of reflection groups
A graph $\mathcal{G}$ is said to be $k$-connected if it contains more than $k$ vertices and remains connected after removing any $k-1$ vertices together with edges incidence to them. An elliptic connection of $(\mathcal{G},\omega)$ is an edge $e$ in $\mathcal{G}$ connecting two nonadjacent vertices on the boundary of a hyperbolic face with $\omega(e)>0$.
@@
Suppose now that $(\mathcal{G},\omega)$ is 2-connected and contains no elliptic connections. For any hyperbolic face $F$ of $\mathcal{G}$, it must bound a Jordan domain and any additional edges connecting vertices of $F$ have weight 00. This means that the cycle of disks corresponding to vertices of $F$ separates $\hat{\mathbb{C}}$ into two parts: one of them is $\Pi_{F}$, and the other may contain additional tangency among the disks (but no overlaps). Arguing similarly as [LLM22, §3.1], the domain of discontinuity $\Omega_{F}$ containing $\Pi_{F}$ is simply connected. ∎
@@
Figure 2.2. Two disk patterns with the same graph but different weights. The one on the right contains an elliptic connection.

### v2

@@ 1. Introduction
[…]
@@
[…]
@@
Let $G$ be a discrete group generated by reflections along finitely many circles in $\hat{\mathbb{C}}$ with connected limit set, and let $\operatorname{QC}(G)$ be the quasiconformal deformation space of $G$. Let $\widetilde{G}\lhd G$ be the index 2 subgroup consisting of orientation preserving elements. Then $\widetilde{G}$ is a geometrically finite Kleinian group. Let $)M\coloneqq\widetilde{G}\backslash(\mathbb{H}^{3}\cup\Omega(\widetilde{G}$)) be the corresponding Kleinian orbifold. Then
@@
We remark that the Bounded Image Theorem does not hold when $M$ is not acylindrical, as otherwise the skinning map may be used to produce a hyperbolic structure on the double $DM$ of $M$ (two copies of $M$ glued along the boundary); see [McM90]. But $DM$ is toroidal, since essential cylinders in $M$ become tori in $DM$. In our setting, a more explicit and detailed discussion can be found in [LZ23, §3].
@@
  • the reflection group $G$ can contain parabolic elements, so $\widetilde{G}\backslash\mathbb{H}^{3}$ can contain cusps, thus the injectivity radius can be 0;
@@
  • for every $N\geq 3$, one can construct a sequence of acylindrical reflection groups with totally geodesic convex core boundary and topological complexity $N$, the depth of whose collar about the convex core boundary goes to 0;
@@ 1.1. Disk pattern and reflection group
Let $\mathcal{G}$ be a connected finite simple plane graph with vertex and edge sets $\mathcal{V}$ and $\mathcal{E}$. Let $\omega:\mathcal{E}\longrightarrow\{\frac{\pi}{n}\colon n\in\mathbb{N}_{\geq 2}\}\cup\{0\}$ be some weight function on its edge set. We call such an edge-weighted graph $(\mathcal{G},\omega)$ a Coxeter graph. A disk pattern with combinatorics $(\mathcal{G},\omega)$ is essentially a collection of disks $\mathcal{P}=\{D_{v},v\in\mathcal{V}\}$ in the Riemann sphere $\hat{\mathbb{C}}$ whose intersection pattern and angles are described by the graph $\mathcal{G}$ and the weight function $\omega$ respectively (see Definitions 2.1 and 2.2 for the subtleties involving a parabolic face).
@@
Hence in our setting, the skinning map can be explicitly defined as follows. Given a disk pattern $\mathcal{P}$ associated to the reflection group, the input of the skinning map is represented by a collection of polygons $\{\Pi_{F,\mathcal{P}}\}_{F}$. For each hyperbolic face $F$, if we configure the disk pattern under an appropriate Möbius transformation so that the polygon $\Pi_{F,\mathcal{P}}$ is the unbounded component of $\hat{\mathbb{C}}-\bigcup_{v\in\mathcal{V}}\overline{D_{v}}$ (i.e. it contains $\infty$), then the corresponding component of the skinning image is represented by the bounded component of $\hat{\mathbb{C}}-\bigcup_{v\in\partial F}\overline{D_{v}}$, which is revealed by removing the disks for vertices in the complement of $F$. This bounded component is another polygon $\Pi_{F,\mathcal{P}}^{-}$ with the same number of sides. See Figure 1.1 for an example illustrating the correspondence between reflection groups and circle patterns, as well as this explicit presentation of the skinning map.
(a) The Coxeter graph, where $n$ on an edge means weight $\pi/n$. Note that there are two quadrilateral hyperbolic faces.
(b) The limit set of a corresponding acylindrical reflection group. It is homeomorphic to a Sierpinski carpet.
(c) A disk pattern $\mathcal{P}$, normalized so that the interstice $\Pi_{F,\mathcal{P}}$ for the face $F$ contains $\infty$. The other interstice $\Pi_{F^{\prime},\mathcal{P}}$ is also visible.
(d) Removing the light gray disks reveals the component of the image of the skinning map for the face $F$, which is also a polygon $\Pi_{F,\mathcal{P}}^{-}$.
Figure $1.$ A reflection group, the corresponding disk pattern, and the skinning image.
@@ 1.2. Discrete extremal lengths / widths
We remark that the uniform lower bound in Equation (1.1) is the key in our argument to obtain a uniform upper bound on the skinning diameter. It is a discrete analogue of the reciprocal condition in [Raj17, NR24] (see §1.4.3 for more discussions).
@@
We remark that as one varies reflection groups with the same topological complexity, the hyperbolic polygons in the skinning image can become degenerate. Our Theorem 1.2, in particular Equation (1.3), states that different hyperbolic polygons in the skinning image must degenerate in the same way. Indeed, the only way to go to infinity in the Teichmüller space of such a hyperbolic polygon is when the hyperbolic distance between some pair of non-adjacent edges goes to zero, by a generalization of [LZ23, Prop. 2.7] (which deals with the case when all angles are zero). Then the polygon becomes long and thin along the direction of the pair of edges, and the extremal width for paths connecting the two edges becomes large. Equation (1.3) implies that for all polygons in the skinning image the corresponding extremal width is comparably large, so they are also long and thin in the same way.
@@ 1.4.2. Circle packings and discrete extremal lengths
Circle packings (or, more generally, disk patterns), their deformation spaces, and rigidity problems have been studied in [RS87, Sch91, He99, RHD07, HL13, HL17]. More generally, these problems have been studied for circle packings on complex projective surfaces [Thu22, KMT03, KMT06, Lam21, BW23]. Circle packings and disk patterns have also long been employed to study geometric structures and their deformation spaces [Bro85, Bro86, FS97]. Recently, [LZ23] has explored the connections between the skinning map, renormalization, and circle packings (see also [LZ25]). Our main result suggests that it may be possible to extend the uniform contraction property of the renormalization operator, as discussed in [LZ23], to cases with non-fixed combinatorics.
@@
Quasi-duality has been applied to prove uniformization theorems of metric surfaces [Raj17, RR19, RRR21, EBPC22, Iko22, MW24]. Notably, in [Raj17], it is shown that a metric space homeomorphic to $\mathbb{R}^{2}$ is quasiconformally homeomorphic to an open domain in $\mathbb{C}$ if and only if it is reciprocal, which in particular requires uniform quasi-duality for all quadrilaterals in the space. This result is generalized in [NR24]. See also a recent exposition by Rajala on quasi-duality and its applications [Raj24].
@@ Acknowledgement
The authors thank Y. Minksy for asking the question and suggesting the connection between uniform diameter bound and circle packings. We also thank the referee for many suggestions that improved the exposition of the paper.
@@ 2.1. Realizable Coxeter graphs
Let $\mathcal{G}$ be a connected finite simple plane graph, and let $\mathcal{V}$ and $\mathcal{E}$ be the set of vertices and edges of $\mathcal{G}$ respectively. Let $\omega:\mathcal{E}\longrightarrow[0,\pi/2]$ be a weight function on the set of edges. We call such an edge-weighted graph a Coxeter graph.
@@
It is easy to see that if $F$ is parabolic, then it is either a triangle or a quadrilateral. It is also easy to see that if $F$ is hyperbolic with $n$ sides, then the weights on the edges add up to $(n-2)\pi$.
For a more uniform presentation, we always assume that in the Coxeter graph $(\mathcal{G},\omega)$, there does not exist a pair of adjacent triangular parabolic faces sharing an edge with weight 0. If there does exist such a pair, it is easy to see that the four edges other than the one shared by the two triangles all have weight $\pi/2$. We can remove the common edge, and combine the two parabolic faces into a single quadrilateral parabolic face. This in fact does not affect the combinatorics of the disk patterns encoded by the graph, see Definitions 2.1 and 2.2, as well as the discussion between them.
@@
$\overline{\mathcal{E}}=\mathcal{E}\cup\bigcup_{F\in\mathcal{F}_{pq}}\{e^{F}_{13},e^{F}_{24}\}.$
The weight function satisfies $\overline{\omega}=\omega$ on $\mathcal{E}$, and $\overline{\omega}\equiv 0$ on $\overline{\mathcal{E}}\backslash\mathcal{E}$.
@@
A disk pattern with Coxeter graph $(\mathcal{G},\omega)$ is a collection of closed round disks $\mathcal{P}\coloneqq\{D_{v},v\in\mathcal{V}\}$ in the Riemann sphere $\hat{\mathbb{C}}$ so that
@@
(Triangulation of a general graph) Suppose now $(\mathcal{G},\omega)$ has at least one hyperbolic face. For each hyperbolic face $F$, add a vertex $v_{F}$ in its interior and connect it to all vertices on $\partial F$. We also assign weight 0 to all the new edges. Note that in this new graph, each hyperbolic face with $n$-sides is divided into $n$ hyperbolic triangles. Now for each new hyperbolic triangle, add an additional vertex and connect it to the three vertices of the triangle. We assign weight $\pi/2$ to these new edges. It is not hard to see that the new Coxeter graph has at least $7$ vertices, and has no hyperbolic face. Conditions (A) and (B) remain true. We can thus apply Step 2 and Step 3 to conclude that the new graph is realizable. Removing the additional disks, we conclude that the original graph is realizable.
@@ 2.2. Reflection groups associated to a Coxeter graph
[…]
Let $\Lambda$ and $\Omega$ be the limit set and domain of discontinuity of $\widetilde{G}$ respectively. Let $\Pi_{F}$ be an interstice and suppose $G_{F}$ is the subgroup of $G$ generated by reflections in the circles corresponding to the vertices $v\in\partial F$. Define $\Omega_{F}=\bigcup_{g\in G_{F}}g\cdot\overline{\Pi_{F}}$.
@@
For Part 1, it suffices to show that for any hyperbolic face $F$, the subgroup $G_{F}$ is nonelementary. Indeed, since $\mathcal{H}_{\mathcal{P}}$ gives a fundamental domain, $G_{F}\cdot\overline{\Pi_{F}}$ tiles $\Omega_{F}$. If $G_{F}$ is elementary, then $\hat{\mathbb{C}}-\Omega_{F}$ consists of at most $2$ points. On the other hand, if we set $\widetilde{G}_{F}$ to be the subgroup of orientation-preserving elements in $G_{F}$, then $\widetilde{G}_{F}\backslash\Omega_{F}$ is hyperbolic, so $\Omega_{F}$ admits a hyperbolic metric as well. This is a contradiction.
@@ 2.3. Moduli spaces and reflection groups
[…]
@@
where $\xi\sim\xi^{\prime}$ if they are conjugates of each other by a Möbius transformation. Note that as a representation of a fixed group $G$, we view the image $\xi(G)$ as marked by the group $G$.
@@
Consider the Kleinian 3-orbifold $M\coloneqq\widetilde{G}\backslash(\mathbb{H}^{3}\cup\Omega)$. Its boundary $\partial M=\bigcup_{F}X_{F}$ has a connected component for each hyperbolic face $F$ of $\mathcal{G}$. Conformally, $X_{F}\cong\widetilde{G}_{F}\backslash\Omega_{F}$. In fact, $X_{F}$ can be constructed as the double of $\Pi_{F}$, with punctures or cone points of order $n$ if the corresponding edges have weight 0 or $\pi/n$. The surface $X_{F}$ has an anti-conformal involution $r=r_{F}$ given by exchanging the two copies of $\Pi_{F}$.
@@ 2.4. Limit sets of reflection groups
A graph $\mathcal{G}$ is said to be $k$-connected if it contains more than $k$ vertices and remains connected after removing any k1k-1 vertices together with edges incident to them. An elliptic connection of $(\mathcal{G},\omega)$ is an edge $e$ in $\mathcal{G}$ connecting two nonadjacent vertices on the boundary of a hyperbolic face with $\omega(e)>0$.
@@
Suppose now that $(\mathcal{G},\omega)$ is 2-connected and contains no elliptic connections. For any hyperbolic face $F$ of $\mathcal{G}$, it must bound a Jordan domain and any additional edges connecting vertices of $F$ have weight 0. This means that the cycle of disks corresponding to vertices of $F$ separates $\hat{\mathbb{C}}$ into two parts: one of them is $\Pi_{F}$, and the other may contain additional tangency among the disks (but no overlaps). Arguing similarly as [LLM22, §3.1], the domain of discontinuity $\Omega_{F}$ containing $\Pi_{F}$ is simply connected. ∎
@@
Figure $2.$ Two disk patterns with the same graph but different weights. The one on the right contains an elliptic connection.